Refer to the information provided in Table 24.1 below to answer the questions that follow. Table 24.1 All Numbers are in $Billion Net Taxes Output Consumption (Incomel Spending Investment Government Spending Spending 800 1,200 1,600 2.000 700 1,000 1,300 1,600 100 100 100 100 200 200 200 200 100 100 100 700 7) Refer to Table 24.1. At an output evel of $1,200 billion, the level of aggregate expenditure is B) $1,200 bilion C) s1 A) $1,000 billion. $1.300 billion D) $1.400 billion 8) Refer to Table 24.1. At an output level of $1 200 billion, there is an unplanned inventory change ot B A) positive $100 billion C) negasive $100 billian l) zero D) posilive $10 billion 9) Refer to Tabie 24.1 At an output level of $2.000 billion, the level of aggregate expenditure b A1 $1,500 billion 80 %1,800 billion. C) $1,900 billion 0) $2.000 billion 0) iD) 1200 $1,200 C) $1,600

Can you solve 8? Also can you show me how we find equilibrium level of output from this table
Answer #1

ANSWER:

8) Solution: negative $100 billion
Explanation: The unplanned inventory at an output level of $1,200 billion is $100 billion computed as: Output - (Consumer spending + Investment spending + Government spending) = 1200-1300 = -100 billion

The equilibrium output level is 1,600 units because at 1,600 units the
Output = Consumer spending + Investment spending + Government spending
1600 = 1300 + 200 +100